OpenStudy (anonymous):

#1: You heat two substances, A and B. Both substances change color. When cooled, substance A returns to its original color, but substance B does not. What most likely happened? A. A physical change occurred in both substances. B. A physical change occurred in substance A, and a chemical change occurred in substance B. C. A chemical change occurred in substance A, and a physical change occurred in substance B. D. A chemical change occurred in both substances. ***my answer : A physical in both is that right??

OpenStudy (anonymous):

#2: How many grams are in a 5 pound sample? Express your answer in scientific notation with two significant figures. (1 pound = 453.592 grams) A. 2.26 x 10^2g B. 2.268 x 10^3g C. 1.1 x 10^4g D. 2.3 x 10^3g **Idk this one.. please help!!

OpenStudy (anonymous):

do u understand these?? is #1 correct?? is it A?? and idk #2.. :(

OpenStudy (anonymous):

I don't think #1 is (A). If substance A returned to its original color, then that was probably just a physical change for A, but substance B stayed the new color after cooling. That makes me think that its color change was more than just physical.

OpenStudy (anonymous):

#2 is just a unit conversion. You can set it up like this \[\large 5 Lbs. = 5 Lb \times \frac{453.392g}{Lb}\]

OpenStudy (anonymous):

wait so for #1, if it isn't A,then what is it?

OpenStudy (anonymous):

and idk how to simplify #2 further :(

OpenStudy (anonymous):

You don't know how to multiply 5 by 453.392?

OpenStudy (anonymous):

\[5\times453,392\]

OpenStudy (anonymous):

1C

OpenStudy (anonymous):

=2,266,960 ????

OpenStudy (anonymous):

wait so #1 is C???

OpenStudy (anonymous):

wich is A

OpenStudy (anonymous):

\[\large 5 \space \cancel{pounds} \times \frac{453.392 \space grams}{1 \space \cancel{pound}} = 5 \times 453.392 \space grams.\]

OpenStudy (anonymous):

It is not (A) for #2, @Jonask

OpenStudy (anonymous):

wait so what??? what was the answer to #1?? I'm confused now :(

OpenStudy (anonymous):

no i was just asking

OpenStudy (anonymous):

A)it is

OpenStudy (anonymous):

I gave you the answer to #1 @iheartfood

OpenStudy (anonymous):

C was the answer to #1??? :)

OpenStudy (anonymous):

Why do you think C? Reread what I wrote about that.

OpenStudy (anonymous):

oh so it's B? B. A physical change occurred in substance A, and a chemical change occurred in substance B. is that the answer to #1??

OpenStudy (anonymous):

@CliffSedge ???

OpenStudy (anonymous):

That's my best guess because A returned to its original color, but B didn't.

OpenStudy (anonymous):

okay yhanks!! what about #2 then???

OpenStudy (anonymous):

Did you multiply the numbers?

OpenStudy (anonymous):

yeah i got this: 2,266,960

OpenStudy (anonymous):

That looks a little large. Make sure you put the decimal point in the right place.

OpenStudy (anonymous):

oh 2,266.960 ???

OpenStudy (anonymous):

@CliffSedge ???

OpenStudy (anonymous):

Right. Now for scientific notation, you want it in the form #.##... ×10^n where n is the number of digits after the first.

OpenStudy (anonymous):

B. 2.268 x 10^3g is this my answer???

OpenStudy (anonymous):

Also note that for numbers that represent measurements and include units, you need to consider significant figures. The weight you were given was 5 pounds. That is only one significant figure, but lets' be generous and assume it's 5.0 pounds for two significant figures.

OpenStudy (anonymous):

A. 2.26 x 10^2g or is it this??

OpenStudy (anonymous):

It has to be ×10^3 because the 2 is in the thousands' place. Also see that 2,266.960 would round off to 2,267 and not 2,268.

OpenStudy (anonymous):

so what is my answer then?? A,B,C, or D??

OpenStudy (anonymous):

As I was saying about the significant figures ("sig. figs"), since you are limited to the precision of your least precise value, you can't have more than two sig. figs. in your answer.

OpenStudy (anonymous):

(This is why I don't like multiple choice questions, it all comes down to guess-and-check or process-of-elimination, and it doesn't teach you how to think through it using the concepts)

OpenStudy (anonymous):

yeah :( so heres what i think now based off of what u said.. is the answer D. 2.3 x 10^3g cuz u can't have more than 2 sig figs right?? is that correct???

OpenStudy (anonymous):

@CliffSedge is that correct???

OpenStudy (anonymous):

@Jonask !! does answer D look right to you>???? :D

OpenStudy (unklerhaukus):

yeah d

OpenStudy (anonymous):

mmk thanks :)